# Graphs that admit a Hamilton path are cup-stackable

### v2

@@ 1. Introduction
In a recent paper, Fay, Hurlbert and Tennant [FHT23] introduced a one-player game on a finite connected graph $G$, which they called cup stacking. Initially, there is a single cup placed on every vertex of $G$. Each move consists of removing all the $r$ cups placed on one vertex $x$ and stacking them onto the cups of a vertex $y$ in distance exactly $r$ from $x$, provided that $y$ already has at least one cup on it. The objective is to eventually move all cups onto a single target vertex $t$, in which case $G$ is called $t$-stackable. If $G$ is $t$-stackable for every vertex $t$, then it is called stackable. Note that the number of vertices with cups on them decreases by 1 with every move, so on a stackable graph the game always ends after exactly $n-1$ moves, where $n$ is the number of vertices of $G$. We remark that stackability is not a monotone property under adding edges. In particular, there are graphs that are stackable, which by adding edges turn into a non-stackable graph.
@@
The paper [FHT23] proves that paths, cycles, two-dimensional grids, the $d$-dimensional hypercube $Q_{d}$ for $d\leq 20$, Kneser graphs $K(n,k)$ for $n\geq 3k-1$, and Johnson graphs with diameter 2 are stackable. It also characterizes which complete partite graphs are stackable. The authors also conjectured that the hypercube $Q_{d}$ is stackable for all $d\geq 1$, and more generally that Cartesian products of paths of arbitrary lengths are stackable. They also asked whether all connected Kneser graphs and generalized Johnson graphs are stackable.
@@ 1.1. Our results
The main contribution of this work is to show that graphs that admit a Hamilton path are stackable (Theorem 1). This considerably enlarges the catalogue of graphs that are known to be stackable. In particular, this yields stackability of $d$-dimensional hypercubes for all $d\geq 1$, Cartesian products of paths of arbitrary lengths, connected Kneser graphs and generalized Johnson graphs (Corollary 2), resolving the problems raised by Fay, Hurlbert and Tennant [FHT23].
@@
Lastly, we construct graphs that are not $t$-stackable for any target vertex $t$, which we call strongly non-stackable. In particular, we obtain strongly non-stackable graphs with arbitrarily large minimum degree and connectivity (Theorems 15 and 16, respectively). We also construct families of graphs that are stackable, but which can be made (strongly) non-stackable by adding edges, showing that stackability is not monotone (Theorems 21 and 22). This shows that our result on stackability of graphs with a Hamilton path is not a trivial consequence of the fact that paths are stackable.
@@ 1.2. Related games and rules
In our version of the game, a move of $r$ cups from a vertex $x$ to another vertex $y$ requires that there is a shortest path $P(x,y)$ of length $r$ from $x$ to $y$ in $G$. This is the reason for the non-monotonicity under adding edges, as adding an edge may change shortest paths. A similar game was studied by Veselovac [Ves22], using slightly different terminology (instead of cups being stacked, he considers frogs that jump on each other), where ‘shortest path $P(x,y)$’ is replaced by ‘path $P(x,y)$’, i.e., cups (or frogs) can move along non-geodesic paths [^1] The English version of the abstract of his thesis says ‘shortest paths’, but this is a mistranslation of the original Croatian version. Definition 1.2 in the thesis clearly requires ‘paths’ only, not ‘shortest paths’., a variant of the game that was also mentioned in [FHT23] for further study. Of course, if $G$ is stackable in the geodesic game, then $G$ is also stackable in the non-geodesic game. Furthermore, if $G$ is a tree both variants are identical, as any path in a tree is also a shortest path. In fact, most results in Veselovac’s thesis are about trees. Specifically, he shows that paths, stars, and so-called starfishs and dandelions are stackable. He also proves that the tree $T$ obtained from two stars with at least 3 rays each by joining their centers with an edge is not $t$-stackable for any vertex $t$. The non-geodesic game behaves monotonically with respect to adding edges, namely if we have a stackable graph, then any graph obtained by adding edges is also stackable. In particular, since the path is stackable, any graph that admits a Hamilton path is also stackable. For the rest of this paper, we do not consider this variant of the game, but instead we focus on the game introduced by Fay, Hurlbert and Tennant [FHT23] where cups have to be moved along geodesic (i.e., shortest) paths.
Mitchell [Mit23] considers yet another variant, where ‘shortest path $P(x,y)$’ is replaced by ‘walk $P(x,y)$’, i.e., vertices and edges used during one move may repeat. As pointed out in [FHT23], other variants are of course possible, such as ‘trail $P(x,y)$’, i.e., vertices may repeat, but not edges.
@@ 2. Graphs that admit a Hamilton path
This substantially enlarges the catalogue of known stackable graphs by all graphs that are known to have a Hamilton path. The following corollary lists only those explicity conjectured or mentioned in [FHT23]. To state the result, we write $P_{\ell}$ for the path with $\ell$ vertices. Furthermore, the Cartesian product $G\operatorname{\Box}H$ of two graphs $G=(V,E)$ and $H=(W,F)$ has the vertex set $V\times W$ and and an edge $((v,w),(v^{\prime},w^{\prime}))$ whenever $v=v^{\prime}$ and $(w,w^{\prime})\in F$ or $(v,v^{\prime})\in E$ and $w=w^{\prime}$. The $d$-dimensional hypercube $Q_{d}$ is the graph $\operatorname{\Box}_{i=1}^{d}P_{2}$. It is easy to show by induction that Cartesian products of paths of arbitrary lengths (see (1) below), in particular $Q_{d}$, admit a Hamilton path. For integers $k\geq 1$ and $n\geq 2k+1$, the Kneser graph $K(n,k)$ has as vertices all $k$-element subsets of $\{1,\ldots,n\}$, and edges between disjoint sets. The generalized Johnson graph $J(n,k,s)$ has as vertices all $k$-element subsets of $\{1,\ldots,n\}$, and an edge between any two sets $A$ and $B$ that satisfy $|A\cap B|=s$. To ensure that the graph is connected, we assume that $s<k$ and $n\geq 2k-s+\mathbf{1}_{[s=0]}$, where $\mathbf{1}_{[s=0]}$ denotes the indicator function that equals 1 if $s=0$ and 0 otherwise. Clearly, we have $J(n,k,0)=K(n,k)$. All these graphs were recently shown to have a Hamilton cycle in [MMN23]. In fact, the Petersen graph $K(5,2)=J(5,2,0)=J(5,3,1)$ does not have a Hamilton cycle, but a Hamilton path.
@@
By Theorem 1, a graph is stackable if it has a Hamilton path. The converse implication, however, is not true in general. The smallest graphs that are stackable but do not admit Hamilton paths have 6 vertices, and they are all shown in Figure 1.
@@ 3.3. Powers of bipartite graphs
For a bipartite graph $G$, we write $\Delta(G)\geq 0$ for the difference in size between the partition classes of $G$. Note that $G^{r}$ is bipartite, and furthermore satisfies $\Delta(G^{r})=\Delta(G)^{r}$. Consequently, if $\Delta(G)\geq 2$, then $\Delta(G^{r})\geq 2^{r}\geq 2$, and so $G^{r}$ does not admit a Hamilton path for any $r\geq 1$. The results below thus yield infinitely many graphs that are stackable, but that do not admit a Hamilton path (in a very strong sense the larger $\Delta(G^{r})$ is; so Theorem 1 does not apply).
@@ 3.4. Powers of trees
This proof is illustrated in Figure 4. For any two vertices $x$ and $y$ in $T$, we write $P(x,y)$ for the path from $x$ to $y$ in $T$. Consider $T$ rooted at one of its leaves $r$. The partition of $T$ into paths is obtained by repeatedly removing a path from this rooted tree as follows: We choose a leaf $x$ at maximum distance from $r$, let $y$ be the vertex on $P(r,x)$ farthest from $r$ that has more than one descendant, and we let $y^{\prime}$ be the descendant of $y$ on $P(r,x)$. The path $P(y,x)$ has length at least $k/2$, otherwise $x$ would have distance less than $k$ from one of the other descendant leaves of $y$. It follows that $P(y^{\prime},x)$ has at least $k/2$ vertices. So we remove $P(y^{\prime},x)$ from $T$ as part of the partition and repeat the process, with the same root $r$, until there is only one path left (this last path has length at least $k$ and therefore at least $k+1$ vertices). As removing $P(y^{\prime},x)$ from the tree does not create additional leaves, this yields the desired partition. ∎
@@
The above proof works in fact also for $k\geq 3$, but the statement of the theorem would be void for $k=3,\ldots,71$, because in those cases $k>k^{3/2}/\sqrt{72}$, and the diameter of any tree is always greater or equal to the spread.
The $s$-subdivision of a graph $G$ is the graph obtained by replacing every edge of $G$ by a path of length $s$.
@@ 5. Stackability is not monotone
Fay, Hurlbert and Tennant [FHT23] completely characterized stackability of complete partite graphs. We need the following special case of their more general result. In this statement, $K_{a,b}$ is the complete bipartite graph with partition classes of size $a$ and $b$.

### v3

@@ 1. Introduction
Motivated by the popular sport of speed stacking, where the goal is to quickly stack and unstack cups in various formations, a recent paper of Fay, Hurlbert and Tennant [FHT24] introduced a one-player game on a finite connected graph $G$, which they called cup stacking. Initially, there is a single cup placed on every vertex of $G$. Each move consists of removing all the $r$ cups placed on one vertex $x$ and stacking them onto the cups of a vertex $y$ in distance exactly $r$ from $x$, provided that $y$ already has at least one cup on it. The objective is to eventually move all cups onto a single target vertex $t$, in which case $G$ is called $t$-stackable. If $G$ is $t$-stackable for every vertex $t$, then it is called stackable. Note that the number of vertices with cups on them decreases by 1 with every move, so on a stackable graph the game always ends after exactly $n-1$ moves, where $n$ is the number of vertices of $G$. We remark that stackability is not a monotone property under adding edges. In particular, there are graphs that are stackable, which by adding edges turn into a non-stackable graph.
@@
The paper [FHT24] proves that paths, cycles, two-dimensional grids, the $d$-dimensional hypercube $Q_{d}$ for $d\leq 20$, Kneser graphs $K(n,k)$ for $n\geq 3k-1$, and Johnson graphs with diameter 2 are stackable. It also characterizes which complete partite graphs are stackable. The authors also conjectured that the hypercube $Q_{d}$ is stackable for all $d\geq 1$, and more generally that Cartesian products of paths of arbitrary lengths are stackable. They also asked whether all connected Kneser graphs and generalized Johnson graphs are stackable.
@@ 1.1. Our results
The main contribution of this work is to show that graphs that admit a Hamilton path are stackable (Theorem 1). This considerably enlarges the catalogue of graphs that are known to be stackable. In particular, this yields stackability of $d$-dimensional hypercubes for all $d\geq 1$, Cartesian products of paths of arbitrary lengths, connected Kneser graphs and generalized Johnson graphs (Corollary 2), resolving the problems raised by Fay, Hurlbert and Tennant [FHT24].
@@
Lastly, we construct graphs that are not $t$-stackable for any target vertex $t$, which we call strongly non-stackable. In particular, we obtain strongly non-stackable graphs with arbitrarily large minimum degree and connectivity (Theorems 15 and 16, respectively). We also construct families of graphs that are stackable, but which can be made (strongly) non-stackable by adding edges, showing that stackability is not monotone (Theorems 21 and 22). This demonstrates that our result on stackability of graphs with a Hamilton path is not a trivial consequence of the fact that paths are stackable.
@@ 1.2. Related games and rules
In our version of the game, a move of $r$ cups from a vertex $x$ to another vertex $y$ requires that there is a shortest path $P(x,y)$ of length $r$ from $x$ to $y$ in $G$. This is the reason for the non-monotonicity under adding edges, as adding an edge may change shortest paths. A similar game was studied by Veselovac [Ves22], using slightly different terminology (instead of cups being stacked, he considers frogs that jump on each other), where ‘shortest path $P(x,y)$’ is replaced by ‘path $P(x,y)$’, i.e., cups (or frogs) can move along non-geodesic paths [^1] The English version of the abstract of his thesis says ‘shortest paths’, but this is a mistranslation of the original Croatian version. Definition 1.2 in the thesis clearly requires ‘paths’ only, not ‘shortest paths’., a variant of the game that was also mentioned in [FHT24] for further study. Of course, if $G$ is stackable in the geodesic game, then $G$ is also stackable in the non-geodesic game. Furthermore, if $G$ is a tree both variants are identical, as any path in a tree is also a shortest path. In fact, most results in Veselovac’s thesis are about trees. Specifically, he shows that paths, stars, and so-called starfishes and dandelions are stackable. He also proves that the tree $T$ obtained from two stars with at least 3 rays each by joining their centers with an edge is not $t$-stackable for any vertex $t$. The non-geodesic game behaves monotonically with respect to adding edges, namely if we have a stackable graph, then any graph obtained by adding edges is also stackable. In particular, since the path is stackable, any graph that admits a Hamilton path is also stackable. For the rest of this paper, we do not consider this variant of the game, but instead we focus on the game introduced by Fay, Hurlbert and Tennant [FHT24] where cups have to be moved along geodesic (i.e., shortest) paths.
Mitchell [Mit23] considers yet another variant, where ‘shortest path $P(x,y)$’ is replaced by ‘walk $P(x,y)$’, i.e., vertices and edges used during one move may repeat. As pointed out in [FHT24], other variants are of course possible, such as ‘trail $P(x,y)$’, i.e., vertices may repeat, but not edges.
@@ 2. Graphs that admit a Hamilton path
This substantially enlarges the catalogue of known stackable graphs by all graphs that are known to have a Hamilton path. The following corollary lists only those explicity conjectured or mentioned in [FHT24]. To state the result, we write $P_{\ell}$ for the path with $\ell$ vertices. Furthermore, the Cartesian product $G\operatorname{\Box}H$ of two graphs $G=(V,E)$ and $H=(W,F)$ has the vertex set $V\times W$ and and an edge $((v,w),(v^{\prime},w^{\prime}))$ whenever $v=v^{\prime}$ and $(w,w^{\prime})\in F$ or $(v,v^{\prime})\in E$ and $w=w^{\prime}$. The $d$-dimensional hypercube $Q_{d}$ is the graph $\operatorname{\Box}_{i=1}^{d}P_{2}$. It is easy to show by induction that Cartesian products of paths of arbitrary lengths, in particular $Q_{d}$, admit a Hamilton path (see (1) below). For integers $k\geq 1$ and $n\geq 2k+1$, the Kneser graph $K(n,k)$ has as vertices all $k$-element subsets of $\{1,\ldots,n\}$, and edges between disjoint sets. The generalized Johnson graph $J(n,k,s)$ has as vertices all $k$-element subsets of $\{1,\ldots,n\}$, and an edge between any two sets $A$ and $B$ that satisfy $|A\cap B|=s$. To ensure that the graph is connected, we assume that $s<k$ and $n\geq 2k-s+\mathbf{1}_{[s=0]}$, where $\mathbf{1}_{[s=0]}$ denotes the indicator function that equals 1 if $s=0$ and 0 otherwise. Clearly, we have $J(n,k,0)=K(n,k)$. Kneser graphs and generalized Johnson graphs were recently shown to have a Hamilton cycle in [MMN23], with the only exception of the Petersen graph $K(5,2)=J(5,2,0)=J(5,3,1)$, which has a Hamilton path, but not cycle.
@@
By Theorem 1, a graph is stackable if it has a Hamilton path. The converse implication, however, is not true in general. With the help of a computer, we found the smallest graphs that are stackable but do not admit Hamilton paths. They have 6 vertices and are all shown in Figure 1.
@@ 3.3. Powers of bipartite graphs
For a connected bipartite graph $G$, we write $\Delta(G)\geq 0$ for the difference in size between the partition classes of $G$. Note that any path in $G$ omits at least $\Delta(G)-1$ many vertices. In particular, if $\Delta(G)\geq 2$, then $G$ has no Hamilton path. Furthermore, note that $G^{r}$ is connected and bipartite as well, and satisfies $\Delta(G^{r})=\Delta(G)^{r}$. Consequently, if $\Delta(G)\geq 2$, then $\Delta(G^{r})\geq 2^{r}$, and so the longest path in $G^{r}$ omits at least $12^{r}-$ 1 vertices, and in particular does not admit a Hamilton path for any $r\geq 1$. The results below thus yield infinitely many graphs that are stackable, but that do not admit a Hamilton path (in a very strong sense that the longest path omits at least $12^{r}-$ 1 many vertices; so Theorem 1 does not apply).
@@ 3.4. Powers of trees
This proof is illustrated in Figure 4. For any two vertices $x$ and $y$ in $T$, we write $P(x,y)$ for the path from $x$ to $y$ in $T$. Consider $T$ rooted at one of its leaves $r$. The partition of $T$ into paths is obtained by repeatedly removing a path from this rooted tree as follows: If the remaining tree is already a path, then the partition consists of the entire path, which has length at least $k$ and therefore at least $k+1$ vertices. Otherwise, we choose a leaf $x$ at maximum distance from $r$, let $y$ be the vertex on $P(r,x)$ farthest from $r$ that has more than one descendant, and we let $y^{\prime}$ be the descendant of $y$ on $P(r,x)$. The path $P(y,x)$ has length at least $k/2$, otherwise $x$ would have distance less than $k$ from one of the other descendant leaves of $y$. It follows that $P(y^{\prime},x)$ has at least $k/2$ vertices. So we remove $P(y^{\prime},x)$ from $T$ as part of the partition and repeat the process, with the same root $r$. As removing $P(y^{\prime},x)$ from the tree does not create additional leaves, this yields the desired partition. ∎
@@
The above proof also works, in fact, for $k\geq 3$, but the statement of the theorem would be void for $k=3,\ldots,71$, because in those cases $k>k^{3/2}/\sqrt{72}$, and the diameter of any tree is always greater or equal to the spread.
The $s$-subdivision of a graph $G$ is the graph obtained by replacing every edge of $G$ by a path of length $s$. The next theorem shows a possible application of Theorem 11. Specifically, the theorem asserts that the third power of any tree $T$ becomes stackable when subdividing the edges of $T$ sufficiently often. This gives a way of constructing more infinite families of stackable powers of trees (recall Remark 7).
@@ 5. Stackability is not monotone
Fay, Hurlbert and Tennant [FHT24] completely characterized stackability of complete partite graphs. We need the following special case of their more general result. In this statement, $K_{a,b}$ is the complete bipartite graph with partition classes of size $a$ and $b$.
@@ 12 $\{22,20,20,18,18,20,20,18,18,20,20,22\}$
From a computational perspective, what is the complexity of deciding whether a given graph $G$ is stackable, or $t$-stackable for some vertex $t$?